Repurchased 205,000 shares of common stock on December 15, Year 3 for $8.00 a share.

Create journal entries and make any necessary changes to the financial statements.

Balance Sheet Statement of Cash HoLS As of1251ear3 For Year Ended 1251Year3 Cash Howfrom Operations Net moome 37,087,771 CurentAsses (51,665,I0) 51,332,0O0 *199,500 (*166,500) 55,554,96953,330,0CO GhangeinAR 55,994,000 55,661,000Chage in Ienry Alowance for Bad Deb5 C333,000) (51,6i5,000)Change in Prepaid Insurance 57,42000 59,324,000thage in Prepaid Rent Prepaid Insurance Prepaid Rent 3199,500 599,000Deprecaion& Aon 3B32,500EN6,I0OChangeinAP 5951,430) 20,539,969 518,315,000GhangeinIncome Tax Payable 51,656,88I Toal CumentASSES Changein erest Payable Loans to uher businesses 266M,I00 2,6KM,00OChage Paid forinterest 32,843,56O52843,560Ghangein Uneamed Revene 55,507,5EO55,507,560Chage in Wages Payable 5199,800 (5166,500) 2,76A28 Net Cash Flowton Opera ons 39,857,000 57,326,000 3A,f52,000 55,328,0I0 55,28,000 Cash Flomfrom estments 518,648,0003B,65B,000Puhase ofLand b2,664,O0) (5B,65B,O0O) (V6,6EO,IIO)Puhase ofEipment Toal PPE tangible Assets Palens, net ToalAsses 322,644,000 511,988,000 NeCash Flowro esimens (512,654 0IO) 5999,0005949,000Cash Howfrom Fnancing 349,640,529 536,804,560 Repayment of Loans Issuance ofNotes Payable Issuance of Bonds Payable Liabiiies and StodholdersEqry Curent Liabies C200,000) Accous Payable terestPayable Income Tax Payable 3,0M4,5O3,96,000NetCash Fow rom Finan 330,000 232880 孔 66.000 $1,198,800 99,000 SEE6,00O 5832,500 5333,0005333,000 5T,595,250 826,500 Net mrease(Decrease) in Cash Cash, Januay 1, Year3 Cash, December 31, Year:3 33,330,0O0 55,554,JGg ages Payable GurentPion ofLoan Payable Toal CumetLiabS Longtem Debt 33,653,000 53,96,000 Net Bonds Payablke Notes Payable 51,561,447 Total Logten Debt Toal Liabiiies 39,324,0I0 55,328,000 514,548,44759,324,000 22,143,697516,150,500 COTIIon s 32,6EO,II0 52,6EO,IOO 51 par, 4,655,000 aied, 266O,000 oustan) $1,998,000 51,998,000 523,633,803 516,746,032 Reained Eamings (5744,972) 5744,972) Tota Sudholders Eqǐy 327,546,831 520,659,0EO TotalLiabiiies and Sactholder549,690,529 536,804,560
Answer #1

Following will be the journal entry for repurchase of common shares of the company :-

General journal debit credit

Treasury stock

Cash

(To 205000 shares repurchased @ $8 per share)

1640000

1640000

Following will be the changes in the financial statements

In balances here

Cash for year 3 will be reduced by $1640000. Thus, Cash = 5554969-1640000 = $3914969.

This, total current assets for year 3 will be 20539969-1640000 = $18899969.

Equity section for year 3 will be as follows :-

Common stock 2660000
Additional paid in capital 1998000
Total paid in capital 4658000
Retained earnings 23633803
Accumulated OCI (744972)
Subtotal 27546831
Less : Treasury stock (205000 shares @$8) (1640000)

Total stockholders equity

$25906831

Balance sheet total for year 3 = $49690529-1640000 = $48050529.

Also there will be change in cash flow statement

Amount paid for repurchase of own share is a neagtive cash flow under financing activities

Thus Net cash Cash flow from financing activities would be $5021969-$1640000 = $3381969.

Following will be the summarized situation of cash flow statement

Net cash flow from operations $9857000
Net cash flow from investment ($12654000)
Net cash flow from financing $3381969
Net increase or (decrease) in cash $584969
Cash , Jan 1, year 3 $3330000
Cash , December 31, year 3 $3914969

We can see that Cash calculate as on December 31, year 3 is same as we have recalculated above.